# DFAMS: Dynamic-flow guided Federated Alignment based Multi-prototype Search

### v1

@@ Abstract
Federated Retrieval (FR) routes queries across multiple external knowledge sources, to mitigate hallucinations of LLMs, when necessary external knowledge is distributed. However, existing methods struggle to retrieve high-quality and relevant documents for ambiguous queries, especially in cross-domain scenarios, which significantly limits their effectiveness in supporting downstream generation tasks. Inspired by dynamic information flow (DIF), we propose DFAMS, a novel framework that leverages DIF to identify latent query intents and construct semantically aligned knowledge partitions for accurate retrieval across heterogeneous sources. Specifically, DFAMS probes the DIF in LLMs by leveraging gradient signals from a few annotated queries and employing Shapley value-based attribution to trace neuron activation paths associated with intent recognition and subdomain boundary detection. Then, DFAMS leverages DIF to train an alignment module via multi-prototype contrastive learning, enabling fine-grained intra-source modeling and inter-source semantic alignment across knowledge bases. Experimental results across five benchmarks show that DFAMS outperforms advanced FR methods by up to 14.37% in knowledge classification accuracy, 5.38% in retrieval recall, and 6.45% in downstream QA accuracy, demonstrating its effectiveness in complex FR scenarios.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) leverages external knowledge documents (Edge et al. 2024; Asai et al. 2023) to effectively enhance the factuality and verifiability of outputs from Large Language Models (LLMs) (OpenAI 2022, 2023; Kaplan et al. 2020; Vu et al. 2024), significantly mitigating issues such as hallucination and knowledge obsolescence (Ji et al. 2023; Cao et al. 2020; Jiang et al. 2024b, a; Asai et al. 2024; Su et al. 2024; Jeong et al. 2024; Baek et al. 2025). However, mainstream RAG approaches typically rely on a single, centralized vector database for knowledge base selection (Kukreja et al. 2024; Bhavnani and Wilson 2009). In reality, knowledge is inherently distributed across multiple heterogeneous data sources — for instance, in medical scenarios, retrieval may need to simultaneously access electronic health records (EHR) (Yuan, Zhou, and Yu 2023), textbooks, and the latest research papers (Zhao et al. 2025).
Figure 1: Since the underlying intent of different users varies, answering the same question may require distinct types of knowledge for a student and a doctor. Such differences in information needs can be reflected in the dynamic information flow within LLMs.
@@
Forcibly aggregating all documents into a unified index not only incurs high retrieval costs but also raises concerns around data sovereignty (Jiang 2024; Shokouhi, Si et al. 2011; Kairouz et al. 2021). To address this, Federated Retrieval (FR) has emerged as a solution, enabling efficient cross-knowledge-base routing decisions that directly and precisely direct queries to the most relevant sources of knowledge (Schölkopf 2019; Guerraoui et al. 2025; Wang et al. 2024c; Ryan et al. 2025; Shojaee et al. 2025).
@@
Most existing FR methods are primarily designed with an emphasis on routing efficiency, privacy preservation, and downstream task integration (Chakraborty, Dahal, and Gupta 2025). However, in real-world scenarios with complex semantics (Clarke et al. 2008), users are more concerned with whether the system can accurately retrieve highly relevant documents to effectively support downstream generation tasks (Shokouhi, Si et al. 2011; Huang and Huang 2024), which does not receive sufficient attention in existing research. And due to limitations in current modeling strategies (Wang et al. 2024c), existing approaches exhibit notable shortcomings in addressing this issue(Guerraoui et al. 2025). On one hand, user queries often suffer from semantic ambiguity or compression, leading to a gap between surface expressions and underlying intent (Yuan et al. 2025). In different contexts, the same question may require different knowledge sources to answer (as illustrate in Figure 1). In such scenarios, user queries often fail to align with the structured and detailed content in the knowledge base, limiting the accuracy and coverage of semantic matching in traditional FR methods (Huang et al. 2021). Although some approaches leverage LLMs for prompt-based query rewriting (Gao et al. 2022) to reduce ambiguity, they often struggle to capture fine-grained semantics due to limited prompt expressiveness, leading to suboptimal performance in semantically complex scenarios. On the other hand, in real-world applications knowledge bases are typically partitioned by data sources, forming multiple structured yet interrelated knowledge subsystems (Wu et al. 2025). The semantic boundaries between these subsystems are often flooded and overlapping. However, some existing methods overlooking underlying semantic connections, which struggle to support cross-source recall and dynamic integration. (Wang et al. 2024c).
Recent research work has shown that when LLMs process tasks of different fields, the contribution of each parameter in the LLM model varies (Dhamdhere, Sundararajan, and Yan 2018; Yu et al. 2018; LeCun, Denker, and Solla 1989). Recent studies on the structural and functional mechanisms of LLMs have revealed that, during inference, LLMs naturally form a Dynamic Information Flow (DIF)—a latent path through which information propagates dynamically across transformer layers, activating neural substructures associated with semantics, knowledge, and reasoning (Zheng et al. ; Yu and Ananiadou 2024; Wang et al. 2024d). These findings suggest that LLMs may already possess an implicit capability to recognize user intent and organize knowledge into latent subdomains when processing complex queries.
@@
Recent studies have shown that LLMs exhibit functional partitioning, where different parameter groups specialize in tasks across diverse domains (Dhamdhere, Sundararajan, and Yan 2018; Yu et al. 2018; LeCun, Denker, and Solla 1989). During inference, these functional partitions associated with semantics, knowledge, and reasoning are selectively activated by the model, giving rise to Dynamic Information Flow (DIF)  (Zheng et al. ; Yu and Ananiadou 2024; Wang et al. 2024d; Stolfo, Belinkov, and Sachan ). Inspired by this line of work, we raise a central research question: Can the dynamic information flow (DIF) within LLMs be explicitly modeled to (1) more accurately identify users’ latent query intents, and (2) structurally segment and dynamically organize overlapping or fuzzy-boundary knowledge subdomains to mitigate semantic misalignment and cross-source retrieval failures in FR settings? As shown in Figure 1, we hypothesize that, when presented with a complex query, the LLM first identifies the user’s intent. If the model determines that the user is likely a doctor with an intent to retrieve clinical studies and guidelines, it will activate the corresponding neurons, forming a reasoning pathway that generates perception signals related to the target knowledge. These signals may be used to retrieve and align content from heterogeneous sources such as PubMed and EHR, integrating cross-subdomain knowledge for downstream generation.
@@
  • We reveal a high-dimensional, information-rich DIF in LLMs that encodes both user intent and subdomain knowledge, enabling more faithful query understanding. To our knowledge, this is the first work to exploit FID for intent-aware, domain-sensitive retrieval modeling.
@@
  • We develop an enhanced FR benchmark dataset to encompass more realistic and diverse query types—ranging from knowledge-free queries, multi-fragment queries within a single source, to cross-source retrieval. The benchmark integrates structured taxonomy, associated documents, user queries, and ground-truth answers, offering a solid foundation for evaluating FR in complex settings.
@@ 2.1 Federated Retrieval
Federated search (Shokouhi, Si et al. 2011), extended into RAG by combining privacy-preserving federated learning (FL) (Zhang et al. 2021) with RAG (Lewis et al. 2020), enables retrieval across decentralized sources without sharing raw data(Chakraborty, Dahal, and Gupta 2025), which has seen widespread adoption in privacy-critical domains such as healthcare (Jiang 2024; Jung, Jeong, and Huh 2025; Xiong et al. 2024), finance, and legal services (Addison et al. 2024). Prior work in federated search mainly targets three aspects: (i) privacy and security through secure retrieval and encryption (Jeon et al. 2021; Peng et al. 2021), (ii) retrieval efficiency via query routing (Wang et al. 2024c), and (iii) integration of FL and RAG tailored to domain-specific tasks (Wang et al. 2024a; Zeng et al. 2024; Shojaee et al. 2025). These advances have proven effective (Zhao 2024; Xu et al. 2022; Wang et al. 2024a; Zeng et al. 2024; Shojaee et al. 2025); however, retrieving high-quality results in complex semantic scenarios remains challenging (Wang et al. 2024c). Existing FR methods, prompt-based and embedding-based, struggle in this setting because queries often misalign with knowledge structures, which reduces dense vector accuracy (Huang et al. 2021), and LLM-based prompt rewriting lacks fine-grained semantic precision (Gao et al. 2022).
@@ 2.2 Neural Information Flow
Recent studies have shown that LLMs, like the human cortex (Arbib 2003; Hawrylycz et al. 2012; Zador 2019; Wang et al. 2024b), exhibit functional partitioning across their architecture (Dhamdhere, Sundararajan, and Yan 2018; Yu et al. 2018; LeCun, Denker, and Solla 1989). Such functional partitions may emerge in the form of attention heads (Zheng et al. 2024; Yin and Steinhardt 2025; Wu et al. 2024), feed-forward networks (Bandarkar et al. 2024; Wendler et al. 2024; Sun et al. 2025), or neurons (Huo et al. 2024; Tang et al. 2024), which are shaped during training and contribute differently across tasks (Dhamdhere, Sundararajan, and Yan 2018; Yu et al. 2018). Building on this partitions, information dynamically flows among these functional modules, forming a Dynamic Information Flow (DIF) (Stolfo, Belinkov, and Sachan ; Yu and Ananiadou 2024) To leverage DIF for downstream tasks, researchers attempt to detect DIF by quantifying the attribution of parameters in the LLM with respect to input query. Quantitative attribution methods have been extensively explored using various techniques, including forward-based methods (Liang et al. ; Todd et al. 2023; Jiang et al. 2025; Dai et al. 2021) and backward-based methods (Feng et al. 2025, 2024) or their combinations (Xu et al. 2024). Among these methods, Shapley value-based approaches (Ghorbani and Zou 2020; Adamczewski, Li, and van Gool 2024) have gained wide adoption. Although there has been extensive research on FIDs in large models, leveraging their powerful capabilities for modeling user intent and knowledge bases remains largely unexplored.
@@ 3.1 Problem Definition
$f_{\text{route}}:x\mapsto\mathbf{w}=[w_{1},w_{2},\dots,w_{I}],\quad w_{j}\in\mathbb{N}_{0},$ (1)
@@
Following RAGRoute (Guerraoui et al. 2025), we combine two types of knowledge: (i) parameterized knowledge $\Theta$ stored in the model weights, and (ii) non-parameterized knowledge $\mathcal{D}=\{\mathcal{K}_{1},\dots,\mathcal{K}_{I}\}$, representing distributed, domain-specific corpora. Given a query $x$, the goal is to generate a reliable response:
$\texttt{Response}\leftarrow\Theta(x,R\mid\mathcal{P}),$ (2)
where $\mathcal{P}$ is a task-specific prompt, and $R$ is the subset of knowledge bases deemed relevant. To handle queries answerable without external retrieval, we include an Others category solved solely by $\Theta$ (Su et al. 2024). During training, supervision is single-source: each $(x,\mathcal{K}_{i},y)$ is paired with a single knowledge base or labeled as Others. At inference, the model generalizes to no-source, single-source, or multi-source retrieval by predicting per-source relevance scores and selecting those above a dynamic threshold $\delta$.
@@ 3.2 Dynamic Information Flow Modeling
To accurately interpret user intent and uncover domain-relevant semantics for routing (C1), we detect Dynamic Information Flow (DIF) capturing which neurons contribute most to domain-sensitive behavior, producing a compact, semantically grounded representation.
The process consists of two steps: (1) constructing a controlled probing dataset to isolate domain-selection behaviors; and (2) identifying key layers and neurons via gradient-based Shapley attribution, and aggregating their activations into a DIF embedding.
@@
Using benchmark training or test sets for attribution often yields spurious signals, as they contain mixed user specific intents. To address this, we construct a dedicated probing dataset:
$\mathcal{D}_{\text{probe}}=\{(x_{i},\mathcal{K}_{i})\}_{i=1}^{n_{\text{probe}}},$
where each query $x_{i}$ is synthesized with fixed instruction-style prompts explicitly designed to elicit domain-selection (e.g., asking the model to identify the most relevant knowledge base). Each label $\mathcal{K}_{i}\in\{1,\dots,I\}$ specifies the correct knowledge base. This dataset is disjoint from $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ to ensure no leakage.
@@
Using $\mathcal{D}_{\text{probe}}$, we estimate the importance of each neuron using Shapley-based attribution. For a transformer block at layer $t$, the feed-forward network (FFN) computes each neuron activation as:
$\theta_{t,j}=\textsc{Act}([h_{t}W_{t1}+b_{t1}]_{j}),$ (3)
where $h_{t}\in\mathbb{R}^{d}$ is the output of the attention sublayer, $W_{t1}\in\mathbb{R}^{d\times 4d}$ and $b_{t1}\in\mathbb{R}^{4d}$ are projection weights and biases, and $\textsc{Act}(\cdot)$ denotes the nonlinearity. For each parameter $\theta_{j}\in\Theta$, we approximate its Shapley value $\phi_{j}$:
@@
where $g_{j}^{(\gamma)}=\partial\mathcal{L}_{\text{SFT}}/\partial\theta_{j}$ is the supervised loss gradient, and $H_{jk}^{(\gamma)}$ is the Hessian capturing second-order interactions. The coefficients $\omega_{jj}^{(j)}$ and $\omega_{jk}^{(\mathcal{S})}$ weight self and pairwise contributions, respectively.
We then: ❶. Select the top-$T$ layers based on aggregated Shapley scores. ❷. Within each selected layer, identify the most informative neuron groups (adjacent units with the highest $\phi_{j}$). ❸. For a query $x$, concatenate the activations of these groups to form the DIF representation:
$\mathbf{z}=\textsc{Concat}\big{(}\{h_{\ell_{i}}^{(g_{i})}\}_{i=1}^{P}\big{)},$ (5)
where $\ell_{i}\in\mathcal{L}_{\text{top}}$ and $g_{i}$ indexes the selected neuron groups. The resulting $\mathbf{z}$ captures both semantic intent and domain cues, and serves as the input for for both inter- and intra-knowledge-base modeling. (Section 3.3).

### v2

@@ Abstract
Federated Retrieval (FR) routes queries across multiple external knowledge sources, to mitigate hallucinations of LLMs, when necessary external knowledge is distributed. However, existing methods struggle to retrieve high-quality and relevant documents for ambiguous queries, especially in cross-domain scenarios, which significantly limits their effectiveness in supporting downstream generation tasks. Inspired by Dynamic Information Flow (DIF), we propose DFAMS, a novel framework that leverages DIF to identify latent query intents and construct semantically aligned knowledge partitions for accurate retrieval across heterogeneous sources. Specifically, DFAMS probes the DIF in LLMs by leveraging gradient signals from a few annotated queries and employing Shapley value-based attribution to trace neuron activation paths associated with intent recognition and subdomain boundary detection. Then, DFAMS leverages DIF to train an alignment module via multi-prototype contrastive learning, enabling fine-grained intra-source modeling and inter-source semantic alignment across knowledge bases. Experimental results across five benchmarks show that DFAMS outperforms advanced FR methods by up to 14.37% in knowledge classification accuracy, 5.38% in retrieval recall, and 6.45% in downstream QA accuracy, demonstrating its effectiveness in complex FR scenarios. Our code are anonymous available at https://anonymous.4open.science/r/DFAMS/
DFAMS: Dynamic-flow guided Federated Alignment based Multi-prototype Search
Zhibang Yang1,2thanks: Equal contribution., Xinke Jiang1,211footnotemark: 1, Rihong Qiu1,2,311footnotemark: 1, Ruiqing Li1,2, Yihang Zhang1, Yue Fang1,2, Yongxin Xu1,3, Hongxin Ding1,2, Xu Chu1,3, Junfeng Zhao1,3thanks: Corresponding authors., Yasha Wang1,2,4 1National Engineering Research Center for Software Engineering, Peking University, Beijing, China 2School of Computer Science, Peking University, Beijing, China 3Key Laboratory of High Confidence Software Technologies, Ministry of Education, Beijing, China 4Peking University Information Technology Institute (Tianjin Binhai), Tianjin, China {yangzb, XinkeJiang, RihongQiu, YueFang}@pku.stu.edu.cn
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) leverages external knowledge documents Edge et al. (2024); Asai et al. (2023) to effectively enhance the factuality and verifiability of outputs from Large Language Models (LLMs) OpenAI (2022, 2023); Vu et al. (2024), significantly mitigating issues such as hallucination and knowledge obsolescence Ji et al. (2023); Cao et al. (2020); Jiang et al. (2024b, a); Asai et al. (2024); Su et al. (2024); Jeong et al. (2024); Baek et al. (2025). However, mainstream RAG approaches typically rely on a single, centralized vector database for knowledge storage and retrieval Kukreja et al. (2024); Bhavnani and Wilson (2009). In reality, knowledge is inherently distributed across multiple heterogeneous data sources — for instance, in medical scenarios, retrieval may need to simultaneously access electronic health records (EHR) Yuan et al. (2023), textbooks, and the latest research papers Zhao et al. (2025).
Figure 1: Hypothesized process of dynamic information flow (DIF) within LLMs for knowledge base selection. When a user asks “Can hypertensive patients use ibuprofen?”, the LLM first infers the latent intent—where a student seeks basic pharmacological understanding, while a doctor requires clinical evidence. The identified intent (e.g., as a doctor) triggers distinct neural and knowledge activations, forming DIF signals that guide retrieval: clinical pathways access ibuprofen records in EHR, whereas conceptual pathways retrieve NSAID-related information from PubMed.
@@
Forcibly aggregating all documents into a unified index not only incurs high retrieval costs but also raises concerns around data sovereignty Jiang (2024); Shokouhi et al. (2011); Kairouz et al. (2021). To address this, Federated Retrieval (FR) has emerged as a solution, enabling efficient cross-knowledge-base routing decisions that directly and precisely direct queries to the most relevant sources of knowledge Schölkopf (2019); Guerraoui et al. (2025); Wang et al. (2024c); Ryan et al. (2025); Shojaee et al. (2025).
@@
Most existing FR methods are primarily designed with an emphasis on routing efficiency, privacy preservation, and downstream task integration Chakraborty et al. (2025). However, in real-world scenarios with complex semantics Clarke et al. (2008), users are more concerned with whether the system can accurately retrieve highly relevant documents to effectively support downstream generation tasks Shokouhi et al. (2011); Huang and Huang (2024), which does not receive sufficient attention in existing research. And due to limitations in current modeling strategies Wang et al. (2024c), existing approaches exhibit notable shortcomings in addressing this issueGuerraoui et al. (2025). On one hand, user queries often suffer from semantic ambiguity or compression, leading to a gap between surface expressions and underlying intent Yuan et al. (2025). In different contexts, the same question may require different knowledge sources to answer (as illustrate in Figure 1). In such scenarios, user queries often fail to align with the structured and detailed content in the knowledge base, limiting the accuracy and coverage of semantic matching in traditional FR methods Huang et al. (2021). Although some approaches leverage LLMs for prompt-based query rewriting Gao et al. (2022) to reduce ambiguity, they often struggle to capture fine-grained semantics due to limited prompt expressiveness, leading to suboptimal performance in semantically complex scenarios. On the other hand, in real-world applications knowledge bases are typically partitioned by data sources, forming multiple structured yet interrelated knowledge subsystems Wu et al. (2025). The semantic boundaries between these subsystems are often flooded and overlapping. However, some existing methods overlooking underlying semantic connections, which struggle to support cross-source recall and dynamic integration. Wang et al. (2024c).
@@
Recent research has shown that when LLMs process tasks of different fields, the contribution of each parameter in the LLM model varies Dhamdhere et al. (2018); Yu et al. (2018); LeCun et al. (1989). Building on this line of work, further investigations into the structural and functional mechanisms of LLMs have revealed that, during inference, LLMs naturally form a Dynamic Information Flow (DIF)—a latent path through which information propagates dynamically across transformer layers, activating neural substructures associated with semantics, knowledge, and reasoning Zheng et al. ; Yu and Ananiadou (2024); Wang et al. (2024d). These findings suggest that LLMs may already possess an implicit capability to recognize user intent and organize knowledge into latent subdomains when processing complex queries. Inspired by this, we raise a central research question: Can DIF within LLMs be explicitly modeled to (1) more accurately identify users’ latent query intents, and (2) structurally segment and dynamically organize overlapping or fuzzy-boundary knowledge subdomains to mitigate semantic misalignment and cross-source retrieval failures in FR settings? As shown in Figure 1, we hypothesize that, when presented with a complex query, the LLM first identifies the user’s intent. If the model determines that the user is likely a doctor with an intent to retrieve clinical studies and guidelines, it will activate the corresponding neurons, forming a reasoning pathway that generates perception signals related to the target knowledge. These signals may be used to retrieve and align content from heterogeneous sources such as PubMed and EHR, integrating cross-subdomain knowledge for downstream generation.
@@
  • We reveal a high-dimensional, information-rich DIF in LLMs that encodes both user intent and subdomain knowledge, enabling more faithful query understanding. To our knowledge, this is the first work to exploit DIF for intent-aware, domain-sensitive retrieval modeling.
@@
  • We develop an enhanced FR benchmark to encompass realistic and diverse query types, ranging from knowledge-free queries, multi-fragment queries within a single source, to cross-source retrieval. The benchmark integrates structured taxonomy, associated documents, user queries, and ground-truth answers, offering a solid foundation for evaluating FR in complex settings.
@@ 2.1 Federated Retrieval
Federated search Shokouhi et al. (2011), extended into RAG by combining privacy-preserving federated learning (FL) Zhang et al. (2021) with RAG Lewis et al. (2020), enables retrieval across decentralized sources without sharing raw dataChakraborty et al. (2025), which has seen widespread adoption in privacy-critical domains such as healthcare Jiang (2024); Jung et al. (2025); Xiong et al. (2024), finance, and legal services Addison et al. (2024). Prior work in federated search mainly targets three aspects: (i) privacy and security through secure retrieval and encryption Jeon et al. (2021); Peng et al. (2021), (ii) retrieval efficiency via query routing Wang et al. (2024c), and (iii) integration of FL and RAG tailored to domain-specific tasks Wang et al. (2024a); Zeng et al. (2024); Shojaee et al. (2025). These advances have proven effective Zhao (2024); Xu et al. (2022); Wang et al. (2024a); Zeng et al. (2024); Shojaee et al. (2025); however, retrieving high-quality results in complex semantic scenarios remains challenging Wang et al. (2024c). Existing FR methods, prompt-based and embedding-based, struggle in this setting because queries often misalign with knowledge structures, which reduces dense vector accuracy Huang et al. (2021), and LLM-based prompt rewriting lacks fine-grained semantic precision Gao et al. (2022).
@@ 2.2 Neural Information Flow
Recent studies have shown that LLMs, like the human cortex Arbib (2003); Hawrylycz et al. (2012); Zador (2019); Wang et al. (2024b), exhibit functional partitioning across their architecture Dhamdhere et al. (2018); Yu et al. (2018); LeCun et al. (1989). Such functional partitions may emerge in the form of attention heads Zheng et al. (2024); Yin and Steinhardt (2025); Wu et al. (2024), feed-forward networks Bandarkar et al. (2024); Wendler et al. (2024); Sun et al. (2025), or neurons Huo et al. (2024); Tang et al. (2024), which are shaped during training and contribute differently across tasks Dhamdhere et al. (2018); Yu et al. (2018). Building on this partitions, information dynamically flows among these functional modules, forming a Dynamic Information Flow (DIF) Stolfo et al. ; Yu and Ananiadou (2024) To leverage DIF for downstream tasks, researchers attempt to detect DIF by quantifying the attribution of parameters in the LLM with respect to input query. Quantitative attribution methods have been extensively explored using various techniques, including forward-based methods Liang et al. ; Todd et al. (2023); Jiang et al. (2025); Dai et al. (2021) and backward-based methods Feng et al. (2025, 2024) or their combinations Xu et al. (2024). Among these methods, Shapley value-based approaches Ghorbani and Zou (2020); Adamczewski et al. (2024) have gained wide adoption. Although there has been extensive research on DIFs in large models, leveraging their powerful capabilities for modeling user intent and knowledge bases remains largely unexplored.
@@ 3.1 Problem Definition
$f_{\text{route}}:x\mapsto\mathbf{w}=[w_{1},w_{2},\dots,w_{I}],\quad w_{j}\in\mathbb{N}_{0},$
@@
Following RAGRoute Guerraoui et al. (2025), we combine two types of knowledge: (i) parameterized knowledge $\Theta$ stored in the model weights, and (ii) non-parameterized knowledge $\mathcal{D}=\{\mathcal{K}_{1},\dots,\mathcal{K}_{I}\}$, representing distributed, domain-specific corpora. Given a query $x$, the goal is to generate a reliable response: $\texttt{Response}\leftarrow\Theta(x,R\mid\mathcal{P})$, where $\mathcal{P}$ is a task-specific prompt, and $R$ is the subset of knowledge bases deemed relevant. To handle queries answerable without external retrieval, we include an Others category solved solely by $\Theta$ Su et al. (2024). During training, supervision is single-source: each $(x,\mathcal{K}_{i},y)$ is paired with a single knowledge base or labeled as Others. At inference, the model generalizes to no-source, single-source, or multi-source retrieval by predicting per-source relevance scores and selecting those above a dynamic threshold $\delta$.
@@ 3.2 Dynamic Information Flow Modeling
To accurately interpret user intent and uncover domain-relevant semantics for routing (C1), we detect DIF capturing which neurons contribute most to domain-sensitive behavior, producing a compact, semantically grounded representation.
The process consists of two steps: (1) constructing a controlled probing dataset to isolate domain-selection behaviors; (2) identifying key layers and neurons via gradient-based Shapley attribution, and aggregating their activations into a DIF embedding.
@@
Using benchmark training or test sets for attribution often yields spurious signals, as they contain mixed user specific intents. To address this, we construct a dedicated probing dataset $\mathcal{D}_{\text{probe}}=\{(x_{i},\mathcal{K}_{i})\}_{i=1}^{n_{\text{probe}}}$, where each query $x_{i}$ is synthesized with fixed instruction-style prompts explicitly designed to elicit domain-selection (e.g., asking the model to identify the most relevant knowledge base). Each label $\mathcal{K}_{i}\in\{1,\dots,I\}$ specifies the correct knowledge base. This dataset is disjoint from $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ to ensure no leakage.
@@
Using $\mathcal{D}_{\text{probe}}$, we estimate the importance of each neuron using Shapley-based attribution. For a transformer block at layer $t$, the feed-forward network (FFN) computes each neuron activation as: $\theta_{t,j}=\textsc{Act}([h_{t}W_{t1}+b_{t1}]_{j})$, where $h_{t}\in\mathbb{R}^{d}$ is the output of the attention sublayer, $W_{t1}\in\mathbb{R}^{d\times 4d}$ and $b_{t1}\in\mathbb{R}^{4d}$ are projection weights and biases, and $\textsc{Act}(\cdot)$ denotes the nonlinearity. For each parameter $\theta_{j}\in\Theta$, its Shapley value $\phi_{j}$ is:
@@
where $g_{j}^{(\gamma)}=\partial\mathcal{L}_{\text{SFT}}/\partial\theta_{j}$ is the supervised loss gradient, and $H_{jk}^{(\gamma)}$ the Hessian for second-order interactions. The coefficients $\omega_{jj}^{(j)}$ and $\omega_{jk}^{(\mathcal{S})}$ weight self and pairwise contributions, respectively.
We then: ❶. Select the top layers $\mathcal{L}_{\text{top}}$ based on aggregated Shapley scores. ❷. Within each selected layer $\ell\in\mathcal{L}_{\text{top}}$, identify the most informative neuron groups $\mathcal{G}_{\ell}$ (adjacent units with the highest $\phi_{j}$). ❸. For a query $x$, collect the activations of these groups and concatenate them to form the DIF representation: $\mathbf{z}(x)=\textsc{Concat}\Big(\big\{h_{\ell}^{(g)}(x)\;\big|\;g\in\mathcal{G}_{\ell}\big\}_{\ell\in\mathcal{L}_{\text{top}}}\Big)$ captures both semantic intent and domain cues, and serves as the input for both inter- and intra-knowledge-base modeling (Section 3.3).